OpenStudy (anonymous):

A kids’ splash pool and an adults’ pool are shaped like similar triangles. The dimensions of the kids’ pool are one-third of the dimensions of the adults’ pool. If the area of the adults’ pool is 189 square meters, what is the area of the kids’ pool?

OpenStudy (anonymous):

63or21

OpenStudy (anonymous):

area of triangle 1=1/2*b*h area of triangle 2=1/2*b/3*h/3=189/9=21

OpenStudy (anonymous):

thnxx